OpenStudy (anonymous):

-3y-6+9=14y-7+3y Pleaseee Help :)

OpenStudy (pfenn1):

\[-3y-6+9=14y-7+3y\] I assume you are trying to solve for y. Do you know how to rearrange the equation to get all y-terms on one side and the rest on the other?

OpenStudy (anonymous):

yeah im solving for y and no...

OpenStudy (pfenn1):

Okay, to get all the y terms on the right side, let's add 3y to both sides of the equation\[\cancel{-3y}-6+9+\cancel{(3y) }=14y-7+3y+(3y)\]resulting in \[3=20y-7\]Now to get the other terms on the left side of the equation we will add 7 to both sides of the equation.\[3+(7)=20y-\cancel{7}+\cancel{(7)}\]resulting in \[10=20y\]Can you solve for y from here?

OpenStudy (anonymous):

yes i can just divide 20 from both sides. Thanks

OpenStudy (anonymous):

answer is 0.5 :)

OpenStudy (pfenn1):

Great.